# 1.3in1.3in1in1in

### v1

@@ 3. On a lagrangian submanifold of $Z$ and singular cubic surfaces
Given a smooth cubic fourfold $Y$, we denote by $Z=Z(Y)$ the associated LLSvS variety. It carries the action of an antisymplectic involution which has been studied in [FMOS23] whose fixed locus consists of two connected components, one isomorphic to the cubic fourfold, which we hence denote $Y$, and a second one which we denote $W$. Goal of this section is to study the less understood variety $W$. In order to do that we introduce a subvariety $P$ of $F\times F$ which is related to $W$ via the Voisin map $\varphi:F\times F\dashrightarrow Z$. A key tool in our strategy is to reduce statements to the study of cubic surfaces $S$ which are given by intersecting the cubic $Y$ with a projective space $\mathbb{P}^{3}$ and consider lines and twisted cubics on $S$.
@@ 3.1. The LLSvS variety $Z$
Given a smooth cubic fourfold $Y$ not containing any plane, Lehn, Lehn, Sorger and van Straten constructed in [LLSvS17] a hyperkähler variety $Z$ of dimension 8, which parametrizes families of twisted cubics on $Y$ and their flat degenerations.
As any twisted cubic has $13t+$ 1 as Hilbert polynomial, the starting point is to consider the compactification $\operatorname{Hilb}^{gtc}(Y)$ of the space of twisted cubics inside the Hilbert scheme $\operatorname{Hilb}^{3t+1}(Y)$. Curves parametrized by this variety are called generalised twisted cubics. The variety $\operatorname{Hilb}^{gtc}(Y)$ is smooth, of dimension 10 and admits a morphism to the Grassmannian $G:=\operatorname{Gr}(4,6)$, which maps any generalised twisted cubic to its linear span. This morphism factors through a ${\mathbb{P}}^{2}$-fibration onto a variety $Z^{\prime}$ which has a generically finite morphism to the Grassmannian. The hyperkähler projective manifold $Z$ is obtained by $Z^{\prime}$ by contracting the divisor parametrising linear systems of curves which are not arithmetically Cohen-Macaulay. It comes with the natural polarisation $\mathcal{L}$ induced by the map to the Grassmannian $G$. It has dimension 8 and the image of the divisor under the contraction is a Lagrangian subvariety isomorphic to the cubic fourfold $Y$.
Remark 3.2.
The construction above of the variety $Z^{\prime}$ is performed relatively over the Grassmannian $G$, so that can be carried out for any cubic surface $S$, thus we shall and will use the variety $Z^{\prime}(S)$ for any cubic surface $S$. It clearly comes equipped with the ${\mathbb{P}}^{2}$-fibration $\mathrm{Hilb}^{gtc}(S)\to Z^{\prime}(S)$.
The hyperkähler manifold $Z$ has a biregular antisymplectic involution $\tau\colon Z\to Z$ as pointed out in [Leh15]. It can be defined as follows: Let $C\subset Y$ be a generalised twisted cubic and let $Q$ be any quadric (surface) in the linear span $\langle C\rangle$, which contains $C$. The intersection $Q\cap Y$ is a sextic which contains $C$, the residual cubic is another generalised twisted cubic $C^{\prime}$. Antisymplectic involutions of K3[n]-type manifolds and this particular example have been studied in [FMOS22, FMOS23].
@@
The fixed locus of the antisymplectic involution has two connected components, one of which is the lagrangian subvariety $Y$. The other component $W$ is of general type, in particular $\omega_{W}=3\mathcal{L}|_{W}$, where $\mathcal{L}$ is the primitive polarization on $Z$.
Voisin constructed a degree 6 rational map $\varphi\colon F\times F\dashrightarrow Z$ in this way [Voi16]. Given a general pair of skew lines $(\ell_{1},\ell_{2})$ and a point $p\in L_{1}$, then the union of $\ell_{1}$ and the quadric, which is residual to $\ell_{1}$ in the intersection $\langle L_{1},\ p\rangle$, is a generalized twisted cubic. The indeterminacy locus, the resolution of it and the branch divisor of its resolution have been studied in [Mur20, Che21, Gio20].
@@
where $\sigma\colon(\ell_{1},\ell_{2})\mapsto(\ell_{2},\ell_{1})$ switches the factors.
@@ 3.4. A variety dominating $W$
$\{(\ell_{1},\ell_{2})\in F\times F:\ell_{i}\text{ are of type I, $\ell_{1}\not=\ell_{2}$ and }\psi(\ell_{1})=\psi(\ell_{2})\}.$
It turns out to have a central rôle in our study: The following intriguing lemma is the starting point which allows us to relate properties of $\psi$ and $\phi$.
@@
Let $(\ell_{1},\ell_{2})\in P$ be a general pair of lines, satisfying the conditions
@@
    and $\psi(\ell_{1})=\psi(\ell_{2})$.
Then the linear span of $L_{1}$ and $L_{2}$ intersects $Y$ in a Cayley cubic surface.
@@
We claim that the linear span $\Lambda:=\langle L_{1},L_{2}\rangle$ intersects $Y$ in a cubic surface with two singular points on $L_{1}$. We denote by $r$ the residual line $\psi(\ell_{1})$, then $R$ and $L_{1}$ are coplanar and we consider the line $T$ in ${\mathbb{P}}(V^{*})$ of hyperplanes which contain $\Lambda$.
@@
If $\psi(\ell_{1})=\psi(\ell_{2})=:r$ then $\varphi(\ell_{1},\ell_{2})=\varphi(\ell_{2},\ell_{1})$.
We report the following theorem which is a key result for our study, but we defer its proof to later, because it relies on some observations conerning Cayley cubics.
@@
The regular map $W\to G$ is birational onto the image, which consists of the projective spaces that intersect $Y$ in a Cayley cubic or a degeneration of it.
In the following we analyze the maps $P\dashrightarrow W\to G$ in greater detail. In order to do so, we describe singular cubic surfaces with focussing on lines on them.
@@ 3.8. Cayley cubics and their degenerations
where ${\widetilde{E}}_{6}$ denotes the cubics with one simple elliptic singularity and $X_{i}$ are non-normal cubic surfaces.
In the following we focus on lines and twisted cubic on singular cubic surfaces to study the maps $\varphi$ and $\psi$. We notice that the map $\psi$ is in general not defined on $Z(S)$ for a singular cubic surface $S$ because a line $L$ need not have a plane tangent at every point. However, if $\pi\subset\langle S\rangle$ is tangent at every point of $L\subset S$ with $\pi\cap S=2L+R$ for a line $R$, we say that the line $R$ is residual, and that $L$ is triple if $R=L$. Generalised twisted cubics on singular cubic surfaces have already been carefully analysed in [LLSvS17, § 2], which we closely follow. Every cubic surface $S$ with at most double rational points has a minimal resolution ${\widetilde{S}}\to S$ with ${\widetilde{S}}$ isomorphic to the blow-up of $\mathbb{P}^{2}$ in six points. The orthogonal $\Gamma:=K_{{\widetilde{S}}}^{\perp}\subset H^{2}({\widetilde{S}},\mathbb{Z})$ is isomorphic to the negative definite root lattice $E_{6}$. We denote by $R_{0}$ the roots in $\Gamma$ which are effective, they correspond to irreducible components of exceptional curves for the resolution. We also denote by $W(R_{0})$ the subgroup of the Weyl group of $E_{6}$ generated by reflections in the effective roots $R_{0}$. The moduli space of generalised twisted cubics on $S$ with reduced structure is then given by a union of projective planes [LLSvS17, Theorem 2.1]:
$\mathrm{Hilb}^{gtc}(S)_{red}=(R/W(R_{0}))\times{\mathbb{P}}^{2}.$
In particular, linear systems of generalised twisted cubics on $S$ are in bijection with the orbits of $R$ under the action of $W(R_{0})$. These correspond to the set $Z^{\prime}(S):=a^{-1}(E)$ where $E\in G$ is the linear span $\langle S\rangle$. Along this bijection families of arithmetically Cohen-Macaulay (aCM) curves correspond to orbits containing an effective root. By the definition of $\tau$ it is clear that it restricts to an action on twisted cubics contained on a cuibic surface, and indeed admits a nice description in this setting: As observed in [Leh15] it is indeed given by the association $\alpha\mapsto-\alpha$ for the roots $\alpha\in R$.
We defer to Appendix 5 details about these singular surfaces, and recall here only the main facts we need to analyse the the Voisin maps.
@@
Any Cayley cubic, i.e. a surface with exactly 4 $A_{1}$ singular points, contains exactly 9 lines whose schematic representation resembles a tetrahedron with vertices at the singular points. We name edges the 6 of them connecting 2 singular points. In particular, if $P_{1},P_{2},P_{3},P_{4}$ are the singular points, we denote $E_{ij}$ the line containing $P_{i}$ and $P_{j}$. Each of the other 3 lines intersect exactly 2 edges, we denote $J_{ij,kl}$ the line joining the edges $E_{ij}$ and $E_{kl}$ for $\{i,j,k,l\}=\{1,2,3,4\}$. See Figure 1(a).
@@
We denote by $K$ the linear system of generalised twisted cubics containing the reducible elements given by the lines through a singular point
$K=[E_{ij}\cup E_{ik}\cup E_{il}]=[E_{ji}\cup E_{jk}\cup E_{jl}]\in Z^{\prime}_{aCM}(S).$
@@
Opposite edges map to $K$ under $\varphi$, that is:
@@
#$\#Z^{\prime}_{aCM}(S)=13\quad\mbox{ and }\quad\#Z^{\prime}(S)^{\tau}=1.$
@@
(of Theorem 3.7) By Lemma 3.5 the general point $(\ell_{1},\ell_{2})$ in $P$ corresponds to opposite edges on a Cayley cubic surface. By Lemma 3.9 its image $\varphi((\ell_{1},\ell_{2}))$ is an aCM twisted cubic fixed by the involution on $Z$, i.e. a point in $W$. In particular, $\varphi$ restricts to a dominant map $\varphi:P\dashrightarrow W$. Finally, the morphism $W\to G$ is 1:1 because by Lemma 3.10 given a projective space $\Lambda\simeq\mathbb{P}^{3}$ which intersects $Y$ in a Cayley cubic $S$, $K$ is the unique linear system of $aCM$ generalised twisted cubics in $S$ which is fixed by $\tau$. ∎
@@
Let $S$ be a cubic surface with a singular point $P_{2}$ of type $A_{3}$, and 2 singular points $P_{0},P_{3}$ of type $A_{1}$. It contains exactly 5 lines as depicted in Figure 1(b): the line $R_{01}$ through $P_{0},P_{1}$; the line $R_{02}$ through $P_{0},P_{2}$; the line $R_{03}$ through $P_{03}$; the line $L_{1}$ containing the only singular point $P_{2}$, and the line $L_{2}$ contained in the smooth locus.
@@
$L_{1}\mbox{ is residual to }R_{23},\qquad L_{1}\mbox{ is residual to }R_{02},$
$L_{2}\mbox{ is residual to }R_{03},\qquad L_{2}\mbox{ is residual to }L_{1}.$
Moreover, there is no plane in $\langle S\rangle\simeq\mathbb{P}^{3}$ tangent to $S$ at every point of $L_{2}$. In particular, there are no triple lines on a surface with $2A_{1}A_{3}$ singularities.
Remark 3.12.
If $S$ is a cubic surface of type $2A_{1}A_{3}$ given by intersecting a smooth cubic fourfold $Y$ with a $\mathbb{P}^{3}$, then one can conclude there are no triple lines on S. Indeed, if all the lines are of type I, Proposition 3.11 can be restated as:
$\psi(R_{23})=\psi(R_{02})=L_{1},\text{ and }\psi(R_{03})=\psi(L_{1})=L_{2}.$
If a line, e.g. $R_{23}$, is of type II one can conclude that ${\widehat{\psi}}(R_{23},\pi)=L_{1}$ where $\pi=\langle R_{23},L_{1}\rangle$.
@@
#$\#Z^{\prime}_{aCM}(S)=5\quad\mbox{ and }\quad\#Z_{aCM}^{\prime}(S)^{\tau}=1.$
@@
$L_{2}\text{ is triple, and }L_{2}\text{ is residual to }L_{1}.$
@@
3.16. Simple elliptic cubic surfaces
Any cubic surface $S$ with a simple elliptic singularity is a cone over a planar elliptic curve $E$. Let $v$ be the vertex. The Fano variety $F(S)$ of $S$ can be naturally identified with the embedded elliptic curve $E$, indeed to any point $P$ of $E$ corresponds the line through $P$ and $v$. The scheme of generalised twisted cubics is isomorphic to $\operatorname{Sym}^{3}(E)$ and the summation map to $E$ is a ${\mathbb{P}}^{2}$-bundle giving the fibration $\operatorname{Hilb}^{gtc}(S)\to Z^{\prime}(S)$ [LLSvS17, Proposition 2.7]. A generalised twisted cubic corresponds to the union of 3 lines, which join the vertex with 3 points on $E$. If these are collinear, then they are the support of a nCM curve with an embedded point at the vertex. Via the isomorphism $F(S)\simeq E$, the residual line can be expressed in terms of the group law of $E$, and the analogue of $P$ for lines on the surface $S$

### v2

@@ 3. On a lagrangian submanifold of $Z$ and singular cubic surfaces
In §3.1, we recall the construction of the hyperkähler variety $Z=Z(Y)$ associated with any smooth cubic fourfold $Y$ not containing a plane, often referred to as the LLSvS variety, together with some of its basic properties. It carries the action of an antisymplectic involution which has been studied in [FMOS22, FMOS23] whose fixed locus consists of two connected components, one isomorphic to the cubic fourfold, which we hence denote $Y$, and a second one which we denote $W$. The goal of this section is to study the less understood variety $W$. In order to do that we introduce in §3.3 we define a subvariety $P$ of $F\times F$ which is related to $W$ via a rational map $\varphi:F\times F\dashrightarrow Z$ introduced by Voisin which we recall in §3.1. A key tool in our strategy is to reduce statements to the study of cubic surfaces $S$ which are given by intersecting the cubic $Y$ with a projective space $\mathbb{P}^{3}$ and consider lines and twisted cubics on $S$. We do so in §3.7. This analysis provides the key tools to prove the main result of the section, Theorem 3.6.
@@ 3.1. The LLSvS variety $Z$
Given a smooth cubic fourfold $Y$ not containing any plane, Lehn, Lehn, Sorger and van Straten constructed in [LLSvS17] a hyperkähler variety $Z$ of dimension 8, which parametrises families of twisted cubics on $Y$ and their flat degenerations. The LLSvS manifold is deformation equivalent to the Hilbert scheme of 4 points on a K3 surface [AL17, Leh18].
As any twisted cubic has $3t+1$ as its Hilbert polynomial, the starting point is to consider the compactification $\operatorname{Hilb}^{gtc}(Y)$ of the space of twisted cubics inside the Hilbert scheme $\operatorname{Hilb}^{3t+1}(Y)$. Curves parametrised by this variety are called generalised twisted cubics. The variety $\operatorname{Hilb}^{gtc}(Y)$ is smooth, of dimension 10 and admits a morphism to the Grassmannian $G:=\operatorname{Gr}(4,6)$, which maps any generalised twisted cubic to its linear span [LLSvS17, Theorem A]. This morphism factors through a ${\mathbb{P}}^{2}$-fibration onto a variety $Z^{\prime}$ followed by a generically finite morphism $a^{\prime}\colon Z^{\prime}\to G$. The hyperkähler projective manifold $Z$ is obtained by $Z^{\prime}$ by contracting the divisor parametrising linear systems of curves which are not arithmetically Cohen-Macaulay. The image of the divisor under the contraction is a Lagrangian subvariety isomorphic to the cubic fourfold $Y$. The variety $Z$ comes with a generically finite rational map $a\colon Z\dashrightarrow G$ and thus a natural polarisation $\mathcal{L}$ induced by the Plücker polarisation of the Grassmannian $G$ [LLSvS17, Theorem B]. The family parametrised by cubic fourfolds form a projective locally complete family of polarised hyperkähler manifolds with polarisation of degree 2 and divisibility 2 [AG23, LPZ23].
The hyperkähler manifold $Z$ has a biregular antisymplectic involution $\tau\colon Z\to Z$ as pointed out in [Leh15]. It can be defined as follows: Let $C\subset Y$ be a generalised twisted cubic and let $Q$ be any quadric (surface) in the linear span $\langle C\rangle$, which contains $C$. The intersection $Q\cap Y$ is a sextic curve consisting of $C$, and the residual generalised twisted cubic $C^{\prime}$. The involution $\tau$ is defined by mapping $C$ to $C^{\prime}$. Antisymplectic involutions of K3[n]-type manifolds and this particular example have been studied in [FMOS22, FMOS23].
@@
The fixed locus of the antisymplectic involution has two connected components, one of which is the lagrangian subvariety $Y$. The other component $W$ is a smooth fourfold of general type, in particular $\omega_{W}=3\mathcal{L}|_{W}$, where $\mathcal{L}$ is the primitive polarisation on $Z$.
Voisin constructed a degree 6 rational map $\varphi\colon F\times F\dashrightarrow Z$ in the following way [Voi16, Theorem 4.8]. Given a general pair of skew lines $(L_{1},L_{2})$ and a point $p\in L_{1}$, then the union of $L_{1}$ and the conic residual to $L_{2}$ in the intersection $\langle L_{2},\ p\rangle\cap Y$ , is a generalised twisted cubic $C$. The map $\varphi$ is defined by the association $(L_{1},L_{2})\mapsto[C]$. The indeterminacy locus, the resolution of it and the branch divisor of its resolution have been studied in [Mur20, Che21, Gio20].
@@
where $\sigma\colon(L_{1},L_{2})\mapsto(L_{2},L_{1})$ switches the factors.
@@ 3.3. A variety dominating $W$
$\{(L_{1},L_{2})\in F\times F:L_{i}\text{ are of type I, $L_{1}\not=L_{2}$ and }\psi(L_{1})=\psi(L_{2})\}.$
The variety $P$ turns out to have a central rôle in our study: The following intriguing lemma is the starting point which allows us to relate properties of $\psi$ and $\phi$.
@@
Let $(L_{1},L_{2})\in P$ be a general pair of lines, satisfying the conditions
@@
    and $\psi(L_{1})=\psi(L_{2})$.
Then the linear span of $L_{1}$ and $L_{2}$ intersects $Y$ in a cubic surface with 4 singular points of type $A_{1}$.
Cubic surfaces with 4 singularities of type $A_{1}$ are called Cayley cubics and we analyse them in §3.7.
@@
We claim that the linear span $\Lambda:=\langle L_{1},L_{2}\rangle$ intersects $Y$ in a cubic surface with two singular points on $L_{1}$. We denote by $R$ the residual line $\psi(L_{1})$, then $R$ and $L_{1}$ are coplanar and we consider the line $T$ in ${\mathbb{P}}(V^{*})$ of hyperplanes which contain $\Lambda$.
@@
If $\psi(L_{1})=\psi(L_{2})=:R$, then $\varphi(L_{1},L_{2})=\varphi(L_{2},L_{1})$.
Proof.
By hypothesis we have $\Lambda_{L_{1}}\cap Y=2L_{1}+R$ and $\Lambda_{L_{2}}\cap Y=2L_{2}+R$. Since $L_{1}$ and $L_{2}$ are skew, each meets $R$ in a single point, say $L_{i}\cap R=\{p_{i}\}$ for $i=1,2$. We compute $\varphi(L_{1},L_{2})$ using the point $p_{1}$, as described in §3.1. Because the construction there applies to a general point of $L_{1}$, we must extend it to this special point $p_{1}$.
The plane $\langle p_{1},L_{2}\rangle$ coincides with $\Lambda_{L_{2}}$, hence its intersection with $Y$ is the cubic curve $2L_{2}+R$. The union $L_{2}\cup R$ is the conic residual to $L_{2}$ in this cubic: This agrees with the definition of residual conic obtained by taking the flat limit. Indeed, set $S:=Y\cap\langle L_{1},L_{2}\rangle$, and consider the morphism $L_{1}\rightarrow\mathrm{Hilb}^{2t+1}(S)$ mapping a point $P$ of $L_{1}$ to the conic in $\langle P,L_{2}\rangle\cap S$ residual to $L_{2}$. Then the limiting residual conic corresponding to $p_{1}$ is precisely $L_{2}\cup R$. Finally, one checks that $L_{1}\cup L_{2}\cup R$ is a point in $\operatorname{Hilb}^{3t+1}(S)$, so that $\varphi(L_{1},L_{2})=L_{1}\cup R\cup L_{2}$. The analogous computation for $\varphi(L_{2},L_{1})$ yields the same conclusion, proving the claim. ∎
We report the following theorem which is a key result for our study, but we defer its proof to later, because it relies on some observations concerning Cayley cubics. We recall that from §3.1 the variety $Z^{\prime}$ admits a morphism $a^{\prime}$ onto the Grassmannian $G=\operatorname{Gr}(4,6)$, which descends to a rational map $a\colon Z\dashrightarrow G$ whose locus of indeterminacy is the connected component of the fixed locus of the involution $\tau$ isomorphic to $Y$.
@@
The regular map $a|_{W}\colon W\to G$ is birational onto the image, which consists of the projective spaces that intersect $Y$ in a Cayley cubic or a degeneration of it.
In the following we analyse the maps $P\dashrightarrow W\to G$ in greater detail. In order to do so, we describe singular cubic surfaces focussing on lines on them.
@@ 3.7. Cayley cubics and their degenerations
(10) $4A_{1}\ \succeq\ 2A_{1}A_{3}\ \succeq\ A_{1}A_{5}\ \succeq\ {\widetilde{E}}_{6}\ \succeq X_{6}\ \succeq X_{7}\ \succeq X_{8}\ \succeq\ X_{9}$
where ${\widetilde{E}}_{6}$ denotes the cubics with one simple elliptic singularity and $X_{i}$ are non-normal cubic surfaces. We say that a cubic surface $S$ is of type $2A_{1}A_{3}$, if $S$ has exactly 2 singularities of type $A_{1}$ and 1 singularity of type $A_{3}$ and analogously for the other singularities.
@@
(11) $\mathrm{Hilb}^{gtc}(S)_{red}=(\Gamma/W(\Gamma_{0}))\times{\mathbb{P}}^{2}.$
In particular, linear systems of generalised twisted cubics on $S$ are in bijection with the orbits of $\Gamma$ under the action of $W(\Gamma_{0})$. Along this bijection families of arithmetically Cohen-Macaulay (aCM) curves correspond to orbits containing an effective root.
Going again through the construction from [LLSvS17] of $Z$, let $Y$ be a cubic fourfold not containing a plane. Mapping a generalised twisted cubic to its linear span defines a morphism from $\mathrm{Hilb}^{gtc}(Y)$ to the Grassmannian $G=\mathrm{Gr}(4,6)$ which factors through a $\mathbb{P}^{2}$-fibration followed by a generically finite morphism $a^{\prime}:Z^{\prime}\to G$. Let $C$ be a twisted cubic and $E$ be its linear span, the intersection $S:=Y\cap E$ is a cubic surface. If $S$ has ADE singularities, the preimage $Z^{\prime}(S):=(a^{\prime})^{-1}(E)$ is in bijection with the set $\Gamma/W(\Gamma_{0})$ in Equation (11) and the points $Z^{\prime}_{aCM}(S)$ corresponding to aCM curves are in bijection with $(\Gamma\setminus\Gamma_{0})/W(\Gamma_{0})$. By its definition the involution $\tau$ restricts to an action on $Z^{\prime}_{aCM}(S)$, and is given by $\alpha\mapsto-\alpha$ for $\alpha\in\Gamma$, see [Leh15].
We defer to Appendix A details about these singular surfaces, and recall here only the main facts we need to analyse the Voisin maps.
@@
Any Cayley cubic, i.e. a cubic surface with exactly 4 $A_{1}$ singular points, contains exactly 9 lines whose schematic representation resembles a tetrahedron with vertices at the singular points. We name edges the 6 of them connecting 2 singular points. In particular, if $P_{0},P_{1},P_{2},P_{3}$ are the singular points, we denote $E_{ij}$ the line containing $P_{i}$ and $P_{j}$. Each of the other 3 lines intersect exactly 2 opposite edges, we denote $J_{ij,kl}$ the line joining the edges $E_{ij}$ and $E_{kl}$ for $\{i,j,k,l\}=\{0,1,2,3\}$. See Figure 1(a).
@@
We denote by $K$ the linear system of generalised twisted cubics containing the reducible elements given by the lines through any singular point $P_{i}$
$K=[E_{ij}\cup E_{ik}\cup E_{il}]\in Z^{\prime}_{aCM}(S).$
for $\{i,j,k,l\}=\{0,1,2,3\}$.
@@
A pair of opposite edges maps to $K$ under $\varphi$, that is:
@@
$\#Z^{\prime}_{aCM}(S)=13\quad\mbox{ and }\quad\#Z_{aCM}^{\prime}(S)^{\tau}=1,$
where $Z_{aCM}^{\prime}(S)^{\tau}$ denotes the locus of fixed points of the involution $\tau$.
@@
(of Theorem 3.6) By Lemma 3.4 the general point $(L_{1},L_{2})$ in $P$ corresponds to opposite edges on a Cayley cubic surface. By Lemma 3.8 its image $\varphi((L_{1},L_{2}))$ is an aCM generalised twisted cubic fixed by the involution on $Z$, i.e. a point in $W$. In particular, $\varphi$ restricts to a dominant map $\varphi:P\dashrightarrow W$. Finally, the morphism $a|_{W}\colon W\to G$ is generically 1:1 onto its image because by Lemma 3.9 given a projective space $\Lambda\simeq\mathbb{P}^{3}$ which intersects $Y$ in a Cayley cubic $S$, $K$ is the unique linear system of $aCM$ generalised twisted cubics in $S$ which is fixed by $\tau$. ∎
@@
Let $S$ be a cubic surface with exactly a singular point $P_{2}$ of type $A_{3}$, and 2 singular points $P_{0},P_{3}$ of type $A_{1}$. It contains exactly 5 lines as depicted in Figure 1(b): the line $R_{02}$ through $P_{0},P_{2}$; the line $R_{23}$ through $P_{2},P_{3}$; the line $R_{03}$ through $P_{0},P_{3}$; the line $L_{1}$ containing the only singular point $P_{2}$, and the line $L_{2}$ contained in the smooth locus.
@@
[…]
Moreover, there is no plane in $\langle S\rangle\simeq\mathbb{P}^{3}$ tangent to $S$ at every point of $L_{2}$.
Proof.
See Appendix A. ∎
@@
$\#Z^{\prime}_{aCM}(S)=5\quad\mbox{ and }\quad\#Z_{aCM}^{\prime}(S)^{\tau}=1.$
@@
[…]
Proof.
See Appendix A. ∎
@@
3.13.1. Simple elliptic cubic surfaces
Any cubic surface $S$ with a simple elliptic singularity is a cone over a planar elliptic curve $E$. Let $v$ be the vertex. The Fano variety $F(S)$ of $S$ can be naturally identified with the embedded elliptic curve $E$, indeed to any point $P$ of $E$ corresponds the line through $P$ and $v$. The scheme of generalised twisted cubics is isomorphic to $\operatorname{Sym}^{3}(E)$ and the summation map to $E$ is a ${\mathbb{P}}^{2}$-bundle giving the fibration $\operatorname{Hilb}^{gtc}(S)\to Z^{\prime}(S)$ [LLSvS17, Proposition 2.7]. An aCM generalised twisted cubic corresponds to the union of 3 lines, which join the vertex with 3 non-collinear points on $E$. If the 3 points are collinear, then they are the support of a non aCM curve with an embedded point at the vertex. Via the isomorphism $F(S)\simeq E$, the residual line can be expressed in terms of the group law of $E$, and the analogue of $P$ for lines on the surface $S$